GlueX TOF Meeting, February 7, 2017

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, February 7, 2017
11:00 am EST
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from the previous meeting
  3. Run Status
    1. Discriminator Thresholds?
    2. TOF Logbook Review
  4. Calibration Status
  5. NIM Paper (Paul)
  6. Action Item Recap



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Thomas Britton, Brad Cannon, Mark Ito (chair), Simon Taylor, Beni Zihlmann

There is a recording of this meeting on the BlueJeans site. Use your JLab credentials.

Lucite Shield Simulation

Thomas has added the Lucite shield to the geometry. He verified its presence using using the HDDS-to-ROOT converter, but it appears not to be showing up when viewed with hdgeant++. Simon will consult on this. Other than this problem, the simulation machinery is ready to go.

Review of minutes from the previous meeting

We went over the minutes from January 24.

  • The broken FADC was replaced by Sascha Somov.
  • New reference plots for the TOF were added Friday.
  • Mark added a check-box for "Hists OK" on the run-by-run info paper form.

Run Status

  • New HV Values. Beni loaded the new values. No voltages are lower than 1200 V.
  • Current draw and scaler rates as a function of beam rate. We need to study this for the new conditions. We are currently running at "medium" current, 100 nA, with 30 kHz trigger rate on a ? micron diamond.
  • Thresholds on the discriminators can be raised from here, and must be for high rate tests, but are acceptable under current conditions.


Beni thought the TOF timing calibrations were OK, but lately discovered several channels where the constants do not make sense. However, there are some good paddles and they show resolution comparable to that obtained in Spring 16 with the higher voltages we had at that time. The problem was due to fits failing. A new set of constants has been produced and went into the CCDB this morning.

New Base Design

Beni has tested a prototype from Vladimir Popov. The results look good, though there is a signification DC offset at high rate.

Replacing PMTs

Beni has had discussions with Tim Whitlatch and Tom Carstens on changing some of the inner tubes after the run, perhaps as many as 32. They are leaning toward a plan where the tubes are replaced in situ (without dismounting the entire TOF planes).